# Thresholds for the distributed surface code in the presence of memory decoherence

### v1

@@ V.2.2 Sensitivity to the link efficiency
Figure 7: Toric surface code error probability thresholds found for $p_{\text{g}}=p_{\text{m}}$, at various values of the coherence times during entanglement generation $T1_{\text{link}}^{\text{n}}(f_{\text{dec}})=0.03f_{\text{dec}}$ seconds and $T2_{\text{link}}^{\text{n}}(f_{\text{dec}})=0.0075f_{\text{dec}}$ seconds. The $f_{\textrm{dec}}$ considered are on the top horizontal axis of the plot. The other simulation parameters are in the third column of Table 1. The corresponding link efficiency is $\eta_{\text{link}}^{*}(f_{\text{dec}})=2f_{\text{dec}}\cdot 10^{2}$. In case of similar performance for the best protocols found in the GHZ optimization, we show the protocol with the lowest $k$ value. This value is printed above the blue markers. Point (*)(*)( * ) shows calculations for a scenario without decoherence. Point ()($\ddagger$)( ‡ ) shows calculations for a scenario without decoherence and with noiseless SWAP gates.
@@
We now verify that, in this regime, similar thresholds can instead be found by varying the link efficiency via the entanglement generation rate. Specifically, we vary the entanglement success probability by adjusting the total photon detection probability $\eta_{\mathrm{ph}}$. For this investigation, which we report in Fig. 8, we use the parameter set in the fourth column of Table 1. This set contains coherence times during entanglement generation that are ten times higher than the state-of-the-art coherence times of Sec. V.1pompiliRealizationMultinodeQuantum2021 We find $p_{\text{g}}=p_{\text{m}}$ thresholds between $p_{\textrm{th}}=0.035(4)$for a photon detection probability corresponding to $\eta_{\textrm{link}}^{*}\approx 4.7\cdot 10^{2}$ and $p_{\textrm{th}}=0.181(4)$for a photon detection probability corresponding to $\eta_{\textrm{link}}^{*}=2\cdot 10^{3}$. At photon detection probability corresponding to $\eta_{\textrm{link}}^{*}\approx 3.6\cdot 10^{2}$ and lower we are not able to find threshold values.
@@ VI Discussion: feasibility of parameter sets below threshold
The total photon detection probability is the product of multiple probabilities (see App. B.1). Present network experiments utilizing NV centers are particularly limited by two of these: the probability of emitting in the zero-phonon-line (ZPL, $\approx 3\%$bernienHeraldedEntanglementSolidstate2013 and the total collection efficiency ($\approx 10-13\%$). hensenLoopholefreeBellInequality2015a ; pompiliRealizationMultinodeQuantum2021 Both values are expected to increase by Purcell enhancement of the NV center emission, for example with the use of optical microcavities rufResonantExcitationPurcell2021 ; riedelDeterministicEnhancementCoherent2017 or nanophotonic devices. liCoherentSpinControl2015 However, even with such devices, the feasibility remains unclear. For microcavities, predicted ZPL emission and collection probabilities are of the order 10 to 46% rufResonantExcitationPurcell2021 ; riedelDeterministicEnhancementCoherent2017 and 49%, rozpedekNeartermQuantumrepeaterExperiments2019 respectively. Moreover, the successful integration of Purcell-enhanced and optically coherent NV centers in nanophotonic devices remains an open research challenge due to the detrimental effects of surface charges. orphal-kobinOpticallyCoherentNitrogenVacancy2023
@@ VII Conclusion
We performed a thorough optimization of GHZ distillation protocols. However, further improvements in other elements of the distributed architecture could partially bridge the gap with the performance of monolithic architectures. For instance, the surface code decoder could model unfinished GHZ distribution rounds as erasure noise and Bell pair distribution could be scheduled dynamically. On top of that, since our software allows for the implementation of general hardware models, further research could focus on analyzing and understanding a broad range of physical systems in the distributed context. In addition to exploring alternative hardware systems, it would be intriguing to implement a more in-depth model of the system’s micro-architecture.
@@ Appendix A Experimental characterization
The noise models and parameter values we use in this paper are based on experimental observations by Bradley et al. Bradley2019 ; bradleyRobustQuantumnetworkMemory2022 and Pompili et al. pompiliRealizationMultinodeQuantum2021 Experiments were performed on NV centers at temperatures of 3.7 and 4 K. Time scales for gates are based on microwave and radio-frequency pulses for single-qubit gates on the electron and ${}^{13}$ qubits, respectively. Time scales for two-qubit gates are based on phase-controlled radiofrequency driving of the carbon spins interleaved with dynamical decoupling sequences on the electron state, following the scheme described by Bradley et al. Taminiau2014 ; Bradley2019
Characterization of the decoherence model and the associated coherence times was achieved by monitoring the drop in probability of detecting Pauli matrix eigenstates at $t>0$ after preparing the state at $t=0$. For the $T_{1}$ relaxation time, the |$\ket{0}$ and |$\ket{1}$ states were used, and the expectation value $\expectationvalue{Z}$ of a measurement in the $Z$-basis was determined after several delay times—the coherence time then directly follows from the observed exponential drop in the expectation value. With the $\ket{+}$ and $\ket{-}$ states and the expectation $\expectationvalue{X}$ for measurement in the $X$-basis, and with |$\ket{+i}$ and |$\ket{-i}$ and $\expectationvalue{Y}$, the $T_{2}$ coherence time of our model could be determined. For these four states, the observed exponential decay $T_{\text{dec}}$ corresponds to the $T_{2}$ time of our model via $1/T_{2}=1/T_{1}-2/T_{\text{dec}}$.
@@
[…]
Nonetheless, in this paper, we assume that the diamond color centers contain a natural abundance of carbon memory qubits (1.1%). Thus we trade off lower coherence times during entanglement generation for faster gates. This choice was made because it is believed that in future systems entanglement success rates are required to be several orders of magnitude higher than current state-of-the-art. In those regimes, fewer entanglement attempts are required and the influence of decoherence during entanglement generation becomes smaller. It is believed that one then benefits more from having the faster operations that samples with natural concentrations of ${}^{13}$ atoms offer.
[…]
Further research into isotopically engineered diamond defect centers is ongoing. This will likely lead to a better understanding of the trade-off between the ${}^{13}$ concentration and the associated operation times and decoherence rates.
@@ B.1 Bell pair state
Here, $I_{0}$ and $I_{1}$ are modified Bessel functions of the zeroth and first order, and $\sigma(\varphi)$ is the standard deviation of the phase instability. We only include phase uncertainty for the single-click protocol, for which a phase instability of ${}^{\circ}$ corresponds to $\lambda=0.984$humphreysDeterministicDeliveryRemote2018
@@
For other systems, the situation might be different. Here, one can design the excitation pulse to induce a $\pi$ transition on the main transition and a full $2\pi$ rotation on the closest unwanted transition. Tiurev et al. created a model that, based on the energy difference between the main transition and the closest unwanted transition, allows one to estimate $p_{\textrm{EE}}$i.e., both the double excitation probability and the probability of exciting this unwanted transition. tiurevFidelityTimebinentangledMultiphoton2021 Their model shows that, typically, the larger the energy difference is between the two transitions, the smaller $p_{\text{EE}}$ becomes.
@@ D.2 Weighted least-squares fitting procedure
To minimize $Q$ with respect to $\bm{\Delta\beta}^{(t+1)}$, one sets $\partial Q/\partial\bm{\Delta\beta}^{(t+1)}$ to zero. This results in an expression that can be solved for $\bm{\Delta\beta}^{(t+1)}$ constalesChapterExperimentalData2017 :

### v2

@@ V.2.2 Sensitivity to the link efficiency
Figure 7: Toric surface code error probability thresholds found for $p_{\text{g}}=p_{\text{m}}$, at various values of the coherence times during entanglement generation $T1_{\text{link}}^{\text{n}}(f_{\text{dec}})=0.03f_{\text{dec}}$ seconds and $T2_{\text{link}}^{\text{n}}(f_{\text{dec}})=0.0075f_{\text{dec}}$ seconds. The $f_{\textrm{dec}}$ considered are on the top horizontal axis of the plot. The other simulation parameters are in the third column of Table 1. The corresponding link efficiency is $\eta_{\text{link}}^{*}(f_{\text{dec}})=2f_{\text{dec}}\cdot 10^{2}$. In case of similar performance for the best protocols found in the GHZ optimization, we show the protocol with the lowest $k$ value. This value is printed above the blue markers. Point ()(*)( ∗ ) shows calculations for a scenario without decoherence. Point ()($\ddagger$)( ‡ ) shows calculations for a scenario without decoherence and with noiseless SWAP gates.
@@
We now verify that, in this regime, similar thresholds can instead be found by varying the link efficiency via the entanglement generation rate. Specifically, we vary the entanglement success probability by adjusting the total photon detection probability $\eta_{\mathrm{ph}}$. For this investigation, which we report in Fig. 8, we use the parameter set in the fourth column of Table 1. This set contains coherence times during entanglement generation that are ten times higher than the state-of-the-art coherence times of Sec. V.1pompiliRealizationMultinodeQuantum2021 We find $p_{\text{g}}=p_{\text{m}}$ thresholds between $p_{\textrm{th}}=0.035(4)$for a photon detection probability corresponding to $\eta_{\textrm{link}}^{*}\approx 4.7\cdot 10^{2}$ and $p_{\textrm{th}}=0.181(4)$for a photon detection probability corresponding to $\eta_{\textrm{link}}^{*}=2\cdot 10^{3}$. At photon detection probability corresponding to $\eta_{\textrm{link}}^{*}\approx 3.6\cdot 10^{2}$ and lower, we are not able to find threshold values.
@@ VI Discussion: feasibility of parameter sets below threshold
The total photon detection probability is the product of multiple probabilities (see App. B.1). Present network experiments utilizing NV centers are particularly limited by two of these: the probability of emitting in the zero-phonon-line (ZPL, $\approx 3\%$bernienHeraldedEntanglementSolidstate2013 and the total collection efficiency ($\approx 10-13\%$). hensenLoopholefreeBellInequality2015a ; pompiliRealizationMultinodeQuantum2021 Both values are expected to increase by Purcell enhancement of the NV center emission, for example with the use of optical microcavities rufResonantExcitationPurcell2021 ; riedelDeterministicEnhancementCoherent2017 or nanophotonic devices. barclayHybridNanocavityResonant2011 ; liCoherentSpinControl2015 However, even with such devices, the feasibility remains unclear. For microcavities, predicted ZPL emission and collection probabilities are of the order 10 to 46% rufResonantExcitationPurcell2021 ; riedelDeterministicEnhancementCoherent2017 and 49%, rozpedekNeartermQuantumrepeaterExperiments2019 respectively. Moreover, the successful integration of Purcell-enhanced and optically coherent NV centers in nanophotonic devices remains an open research challenge due to the detrimental effects of surface charges. orphal-kobinOpticallyCoherentNitrogenVacancy2023
@@ VII Conclusion
We performed a thorough optimization of GHZ distillation protocols. However, further improvements in other elements of the distributed architecture could partially bridge the gap with the performance of monolithic architectures. For instance, the surface code decoder could model unfinished GHZ distribution rounds as erasure noise. The conversion of binary trees to protocol recipes can be optimized and Bell pair distribution could be scheduled dynamically. On top of that, since our software allows for the implementation of general hardware models, further research could focus on analyzing and understanding a broad range of physical systems in the distributed context. In addition to exploring alternative hardware systems, it would be intriguing to implement a more in-depth model of the system’s micro-architecture.
@@ Appendix A Experimental characterization
The noise models and parameter values we use in this paper are based on experimental observations by Bradley et al. Bradley2019 ; bradleyRobustQuantumnetworkMemory2022 and Pompili et al. pompiliRealizationMultinodeQuantum2021 Experiments were performed on NV centers at temperatures of 3.7 and 4 K. Time scales for gates are based on microwave and radio-frequency pulses for single-qubit gates on the electron and 13C qubits, respectively. Time scales for two-qubit gates are based on phase-controlled radiofrequency driving of the carbon spins interleaved with dynamical decoupling sequences on the electron state, following the scheme described by Bradley et al. Taminiau2014 ; Bradley2019
Characterization of the decoherence model and the associated coherence times was achieved by monitoring the drop in probability of detecting Pauli matrix eigenstates at $t>0$ after preparing the state at $t=0$. For the $T_{1}$ relaxation time, the |$\ket{0}$ and |$\ket{1}$ states were used, and the expectation value $\expectationvalue{Z}$ of a measurement in the Pauli-$Z$ basis was determined after several delay times—the coherence time then directly follows from the observed exponential drop in the expectation value. With the $\ket{+}$ and $\ket{-}$ states and the expectation $\expectationvalue{X}$ for measurement in the Pauli-$X$ basis, and with |$\ket{+i}$ and |$\ket{-i}$ and $\expectationvalue{Y}$, the $T_{2}$ coherence time of our model could be determined. For these four states, the observed exponential decay $T_{\text{dec}}$ corresponds to the $T_{2}$ time of our model via $1/T_{2}=1/T_{1}-2/T_{\text{dec}}$.
The full model is based on the so-called NV- state, a spin-1 electron qubit with spin projection quantum number $m_{s}\in\{-1,0,1\}$. Stochastic ionization can convert the NV- state to the NV0 state with $m_{s}\in\{-1/2,1/2\}$. Under the present understanding, this spin state is no longer usable as a qubit due to a fast orbital relaxation process. baierOrbitalSpinDynamics2020 Since the electron-spin state is used to control the 13C spins, ionization accordingly dephases the 13C states. As such, the coherence times of the NV center memory are currently limited by these ionization effects. Ref. bradleyRobustQuantumnetworkMemory2022, proposes a method to mitigate ionization-induced memory dephasing by actively recharging the NV0 state. They show ionization and recharging can be performed with minimal loss in fidelity in the state of a 13C spin in an isotopically purified device. This marks an important avenue for future research. In our model, we do not specifically include ionization, but simply absorb its influence in the 13C coherence times.
Furthermore, Ref. bradleyRobustQuantumnetworkMemory2022, showed that isotopically engineered nitrogen-vacancy devices with a reduced 13C memory qubit concentration (0.01%) are able to store a quantum state over 105 entanglement attempt repetitions. Compared to samples with natural 13C abundance (1.1%), the memory qubits have a weaker coupling to their color centers. While this increases coherence times during entanglement generation by several orders of magnitude, it also leads to longer time scales for carrying out gates on the memory qubits and gates between the communication and memory qubits. In natural abundance devices, a quantum state can typically be stored over $10^{3}$ entanglement attempt repetitions, pompiliRealizationMultinodeQuantum2021 while demonstrated single qubit 13C gates are typically $\approx 13$ times faster and two-qubit gates are typically $\approx 50$ times faster than the isotopically purified samples.
Nonetheless, in this paper, we assume that the diamond color centers contain a natural abundance of carbon memory qubits (1.1%). Thus we trade off lower coherence times during entanglement generation for faster gates. This choice was made because it is believed that in future systems entanglement success rates are required to be several orders of magnitude higher than current state-of-the-art. In those regimes, fewer entanglement attempts are required and the influence of decoherence during entanglement generation becomes smaller. It is believed that one then benefits more from having the faster operations that samples with natural concentrations of 13C atoms offer.
@@
Further research into isotopically engineered diamond defect centers is ongoing. This will likely lead to a better understanding of the trade-off between the 13C concentration and the associated operation times and decoherence rates.
@@ B.1 Bell pair state
Here, $I_{0}$ and $I_{1}$ are modified Bessel functions of the zeroth and first order, and $\sigma(\varphi)$ is the standard deviation of the phase instability. We only include phase uncertainty for the single-click protocol, for which a phase instability of 14.3 corresponds to $\lambda=0.984$humphreysDeterministicDeliveryRemote2018
@@
For other systems, the situation might be different. Here, one can design the excitation pulse to induce a $\pi$ transition on the main transition and a full $2\pi$ rotation on the closest unwanted transition. Tiurev et al. created a model that, based on the energy difference between the main transition and the closest unwanted transition, allows one to estimate $p_{\textrm{EE}}$i.e., both the double excitation probability and the probability of exciting this unwanted transition. tiurevFidelityTimebinentangledMultiphoton2021 ; mirambellFidelityCharacterizationSpinphoton2019 Their model shows that, typically, the larger the energy difference is between the two transitions, the smaller $p_{\text{EE}}$ becomes.
@@ D.2 Weighted least-squares fitting procedure
To minimize $Q$ with respect to $\bm{\Delta\beta}^{(t+1)}$, one sets $\partial Q/\partial\bm{\Delta\beta}^{(t+1)}$ to zero. This results in an expression that can be solved for $\bm{\Delta\beta}^{(t+1)}$constalesChapterExperimentalData2017